### v1

Abstract
This paper investigates the hull codes of free linear codes over a non-unital ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$. Initially, we examine the residue and torsion codes of various hulls of $E$-linear codes and obtain an explicit form of the generator matrix of the hull of a free $E$-linear code. Then, we propose four build-up construction methods to construct codes with a larger length and hull-rank from codes with a smaller length and hull-rank. Some illustrative examples are also given to support our build-up construction methods. Subsequently, we study the permutation equivalence of two free $E$-linear codes and discuss the hull-variation problem. As an application, we classify optimal free $E$-linear codes for lengths up to $8$.
journal:
1 Introduction
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements. A subspace $C$ of $\mathbb{F}_{q}^{n}$ refers to an $\mathbb{F}_{q}$-linear code of length $n$. Under the Euclidean inner product, the dual of the code $C$ is the collection of all vectors of $\mathbb{F}_{q}^{n}$ that are orthogonal to $C$ and denoted by $C^{\perp}$. The hull of the code $C$ is defined as $Hull(C)=C\cap C^{\perp}$. In fact, the concept of the hull of an $\mathbb{F}_{q}$-linear code was introduced in 1990 by Assmus and Key Assmus to investigate certain properties of finite projective planes. Recall that two codes $C$ and $D$ are permutation-equivalent if $C$ can be derived from $D$ by applying a suitable coordinate permutation. Determining whether two linear codes are permutation-equivalent is a significant problem in coding theory. Recently, hulls have attracted intense research interest due to their various applications. In particular, the dimension of the hull directly influences the complexity of algorithms that check the permutation equivalence or determine the automorphism group of a linear code Leon ; Send1997 ; Send2000 . Additionally, hulls can be utilized to construct quantum error-correcting codes Dougherty ; Guenda . These facts highlight the necessity of studying hull codes. Over the years, researchers have investigated the hulls of various codes over finite fields or finite rings with unity. However, there is no work available in the literature on the hull of codes over non-unital rings. Thus, the study of hulls over non-unital rings is a significant problem, and this is a first attempt towards this direction.
Recall that an $\mathbb{F}_{q}$-linear code $C$ is referred to as a linear complementary dual (LCD), self-orthogonal (SO) and self-dual (SD) code if it satisfies $\cap C^{\perp}=\{0\}, CC$$\subseteq C^{\perp}$ and $C=C^{\perp}$, respectively. We see that, if $Hull(C)$ is trivial, it is LCD, if $Hull(C)=C$, it is SO, and if $Hull(C)=C=C^{\perp}$, it is SD. Thus, the hull code can be seen as a basic generalization of LCD, SO and SD codes. The concept of LCD codes was introduced by Massey in Massey92 , where he also proved the existence of asymptotically good LCD codes. In recent years, these codes have attracted considerable interest because of their use in countering Side-Channel Attacks Carl16 . Using the concept of hull dimension spectra, Sendrier proved in 2004 that LCD codes attain the Gilbert-Varshamov bound Send04 . Subsequently, Islam et al. Islam2022 demonstrated a method to construct both quantum and LCD codes leveraging cyclic codes over non-chain rings. Further, double circulant LCD and SD codes were investigated in Prakash . On the other hand, research on SO and SD codes has seen significant growth due to their use in different areas such as quantum stabilizer codes Calderbank ; Steane , modular forms Minjia , lattice theory Bannai ; Conway ; Harada , and combinatorial $t$-design theory Bachoc .
Recently, there has been a significant shift towards considering non-unital rings to study different types of codes. In 2022, Shi et al. Shi21 were the first to consider a non-unitary ring with $4$ elements to investigate LCD codes. Following this, Kim and Roe Kim2022a investigated quasi-self-dual (QSD) codes over a non-unitary ring with 4 elements. Note that a QSD code $C$ of length $n$ over a finite ring $R$ is an SO code with $|C|=|R|^{n/2}$. Subsequently, authors in Kushwaha1 extended QSD and LCD codes over a non-unitary ring with $9$ elements. Meanwhile, Deb et al. Deb worked on classifying certain codes over a non-unitary ring of order $4$. This was followed by the work of Kushwaha et al. Kushwaha2 , who classified MDS and almost MDS LCD and SD codes over a non-unital ring $E_{p}$. On the other hand, the build-up construction method is a powerful technique to construct codes with a larger length from codes with a smaller length. This technique was utilized in Alah22 and Kim2022a for the classification of Type IV and QSD codes over non-unital rings for some smaller lengths. Further, binary linear codes with various hull dimensions were classified in Li using build-up construction techniques. In 2020, Liu and Pan Liu studied the hull-variation problem for $\mathbb{F}_{q}$-linear codes, while LCD and $\mathbb{F}_{q}$-linear codes with hull-dimension $1$ were investigated in Wang . Motivated by these works, we study the hulls of free codes over the non-unital ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$. To achieve this, we explore the various hulls of the $E$-linear codes and identify their residue and torsion codes. Then, we produce the form of the generator matrix of the hull code of a free $E$-linear code and also determine its hull-rank. Subsequently, we present four build-up construction methods that utilize free $E$-linear codes with a smaller length and hull-rank to construct free $E$-linear codes with a larger length and hull-rank. We also give some supporting examples of codes constructed by using these build-up construction methods. In addition, we investigate the permutation equivalence of $E$-linear codes and discuss the hull-variation problem for free $E$-linear codes. We conclude our work by classifying optimal free $E$-linear codes for lengths up to $8$.
We organize this work in the following manner. Section 2 provides some preliminaries and studies various hulls of an $E$-linear code. It also deals with the generator matrix of the hull code of a free $E$-linear code and its hull-rank. In Section $3$, the four build-up construction methods are presented. Section $4$ investigates the hull-variation problem for free $E$-linear and classifies the optimal free $E$-linear codes for lengths up to $8$. We conclude our work in Section $5$.
2 Hulls over the ring $E$
In Fine93 , Fine has classified rings of order $4$ up to isomorphism. There are only two noncommutative non-unitay rings with $4$ elements, namely, the ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$ and its opposite ring. This paper considers the ring $E$ as a code alphabet to study the hull of free linear codes over $E$. One can derive similar results over the opposite ring. The characteristic of the ring $E$ is $2$ and it has $4$ elements $\{~i\kappa+j\tau~|~0\leq i,j<2~\}$. We denote $\kappa+\tau=\zeta$ and give the multiplication table of the ring $E$ in Table 1 to characterize its structure.
Table 1: Multiplication table of $E$.
$\cdot$ 0 $\kappa$ $\tau$ $\zeta$
0 0 0 0 0
$\kappa$ 0 $\kappa$ $\kappa$ 0
$\tau$ 0 $\tau$ $\tau$ 0
$\zeta$ 0 $\zeta$ $\zeta$ 0
The multiplication table shows that the ring $E$ is non-unitary and has a unique maximal ideal $J=\{0,\zeta\}$. Hence, $E$ is a local ring. Therefore, its residue field is given by $E/J=\mathbb{F}_{2}$. Further, $x\kappa=x\tau=x$ for all $x\in E$. Also, every element $x\in E$ has a $\zeta$-adic decomposition as follows:
$x=u\kappa+v\zeta~~\text{where}~u,v\in\mathbb{F}_{2}.$
Now, for all $x\in E$ and $v\in\mathbb{F}_{2}$, an action of $\mathbb{F}_{2}$ on $E$ can be defined as $xv=vx$. We see that this action is distributive, i.e., $x(u\oplus_{\mathbb{F}_{2}}v)=xu+xv=ux+vx$ for all $x\in E$ and $u,v\in\mathbb{F}_{2}$. Now, if an element of $E$ is written in the form of $\zeta$-adic decomposition, then the map of reduction modulo $J$ is a map $\pi:E\rightarrow E/J=\mathbb{F}_{2}$ defined by
$\pi(x)=\pi(u\kappa+v\zeta)=u.$
There is a natural extension of this map from $E^{n}$ to $\mathbb{F}_{2}^{n}$.
Remark 1
In this paper, we fix the notations $\kappa$ and $\tau$ for the generators of the ring $E$ that satisfy the conditions $2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau$. Moreover, $\zeta=\kappa+\tau.$
Definition 1(Linear code)
An $E$-linear code of length $n$ refers to a left $E$-submodule of $E^{n}$.
Definition 2(Minimum distance)
The number of differing components between two codewords $w$ and $z$ is called the distance between them and is denoted by $d(w,z)$. Then, the minimum distance of the code $C$ is
$d(C)=\text{min}\{d(w,z)~|~w,z\in C,w\neq z\}.$
Here, given an $E$-linear code $C$ of length $n$, the following two $\mathbb{F}_{2}$-linear codes related to $C$ are defined.
  1. 1.
    Residue code: For an $E$-linear code $C$, $\pi(C)$ is called its residue code and denoted by $C_{Res}$, i.e.,
    $C_{Res}=\{\pi(z)~|~z\in C\}.$
  2. 2.
    Torsion code: For an $E$-linear code $C$, its torsion code is an $\mathbb{F}_{2}$-linear code given by
    $C_{Tor}=\{v\in\mathbb{F}_{2}^{n}~|~v\zeta\in C\}.$
Now, we define the inner product of any vectors $w=(w_{1},w_{2},\ldots,w_{n})$, and $z=(z_{1},z_{2},\ldots,z_{n})$ in $E^{n}$ as
$\langle w,z\rangle=\sum_{j=1}^{n}w_{j}z_{j}.$
Additionally, the duals of an $E$-linear code $C$ are introduced below, defined under the inner product given above.
  1. 1.
    Left dual: The $E$-linear code given by
    $C^{\perp_{L}}=\{\boldsymbol{z}\in E^{n}~|~\langle\boldsymbol{z},\boldsymbol{w}\rangle=0,\forall~\boldsymbol{w}\in C\},$
    is said to be the left dual of the code $C$.
  2. 2.
    Right dual: The $E$-linear code given by
    $C^{\perp_{R}}=\{\boldsymbol{z}\in E^{n}~|~\langle\boldsymbol{w},\boldsymbol{z}\rangle=0,\forall~\boldsymbol{w}\in C\},$
    is defined as the right dual of the code $C$.
  3. 3.
    Dual: The intersection of left and right duals of an $E$-linear code $C$ is defined as the dual of the code $C$ and denoted by $C^{\perp}$, i.e., $C^{\perp}=C^{\perp_{L}}\cap C^{\perp_{R}}$.
Definition 3(Left hull, right hull and hull codes)
For an $E$-linear code $C$, its left hull, right hull and hull codes are defined as $LHull(C)=C\cap C^{\perp_{L}}$, $RHull(C)=C\cap C^{\perp_{R}}$ and $Hull(C)=C\cap C^{\perp}$, respectively.
Definition 4(Generating set, generator matrix and parity-check matrix)
For an $E$-linear code $C$, let $X=\{x_{1},x_{2},\ldots,x_{k}\}\subset C$. Then the set
$\langle X\rangle_{E}=\{a_{1}x_{1}+a_{2}x_{2}+\cdots+a_{k}x_{k}~|~a_{i}\in E,~1\leq i\leq k\},$
is called the (left) $E$-span of $X$. Next, the set given by
$\langle X\rangle_{\mathbb{F}_{2}}=\{v_{1}x_{1}+v_{2}x_{2}+\cdots+v_{k}x_{k}~|~v_{i}\in\mathbb{F}_{2},~1\leq i\leq k\},$
is called the additive span of the set $X$. Since the ring $E$ is non-unitary, the additive span of the set $X$ is not always contained in the $E$-span of $X$. Now, if the set $X=\{x_{1},x_{2},\ldots,x_{k}\}$ satisfies
$\langle X\rangle_{E}\cup\langle X\rangle_{\mathbb{F}_{2}}=C,$
it is termed as a generating set for the code $C$. Next, if $X=\{x_{1},x_{2},\ldots,x_{k}\}\subset C$ is a generating set of an $E$-linear code $C$ of length $n$, then a $k\times n$ matrix $G_{E}$ whose rows are $x_{1},x_{2},\ldots,x_{k}$ and $\langle G\rangle_{E}=\langle X\rangle_{E}\cup\langle X\rangle_{\mathbb{F}_{2}}$ refers to a generator matrix of $C$. Further, a generator matrix of $C^{\perp}$ is termed the parity-check matrix of $C$.
Definition 5(Free code)
If an $E$-linear code $C$ can be decomposes as
$C=E\oplus E\oplus\cdots\oplus E,$
where each component $E$ is generated by some $z_{i}\in E$ as an $E$-module, it is called free. Note that the freeness of the $E$-linear code $C$ is equivalent to the condition $C_{Res}=C_{Tor}$.
Definition 6(Rank of a free code)
For a free $E$-linear code $C$, its rank, denoted by $rank(C)$, is the size of its minimal generating set.
We now state a few basic results derived from Lemma $1$, Theorem $9$ and Theorem $10$ of Alah23 .
Theorem 2.1
If $C$ is an $E$-linear code of length $n$, then
  1. 1.
    $\kappa C_{Res}\subseteq C$ and $\zeta C_{Tor}\subseteq C$,
  2. 2.
    $(C^{\perp_{L}})_{Res}=(C_{Res})^{\perp}=(C^{\perp_{L}})_{Tor}$,
  3. 3.
    $(C^{\perp_{R}})_{Res}=(C_{Tor})^{\perp}$ and $(C^{\perp_{R}})_{Tor}=\mathbb{F}_{2}^{n}$,
  4. 4.
    $(C^{\perp})_{Res}=(C_{Tor})^{\perp}$ and $(C^{\perp})_{Tor}=(C_{Res})^{\perp}$.
Now, we have the following result to check the freeness of the three duals of an $E$-linear code.
Theorem 2.2
If $C$ is an $E$-linear code, then $C^{\perp_{L}}$ is always free but $C^{\perp_{R}}$ is never free. Further, $C^{\perp}$ is free if and only if $C$ is free.
Proof
Freeness of $C^{\perp_{L}}$ and non-freeness of $C^{\perp_{R}}$ follow the Lemma $5$ of Shi21 . To prove the necessity of the second result, i.e., for the free $E$-linear code $C$, the result (4) of Theorem 2.1 implies that
$(C^{\perp})_{Res}=(C_{Res})^{\perp}=(C^{\perp})_{Tor}.$
Therefore, $C^{\perp}$ is free. On the other hand, assume $C^{\perp}$ is a free $E$-linear code. Since $(C^{\perp})_{Res}=(C^{\perp})_{Tor}$, again (4) of Theorem 2.1 implies that $(C_{Res})^{\perp}=(C_{Tor})^{\perp}$. Thus, $C_{Res}=C_{Tor}$, and so $C$ is free.
The following result calculates the residue and torsion codes of the three hull codes of an $E$-linear code.
Theorem 2.3
Let $C$ be an $E$-linear code. Then
  1. 1.
    $(LHull(C))_{Res}=C_{Res}\cap(C_{Res})^{\perp}$ and $(LHull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}$,
  2. 2.
    $(RHull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp}$ and $(RHull(C))_{Tor}=C_{Tor}$,
  3. 3.
    $(Hull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp}$ and $(Hull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}$.
Proof
  1. 1.
    Let $u\in(LHull(C))_{Res}$. Then, by (1) of Theorem 2.1, $\kappa u\in C\cap C^{\perp_{L}}$ which implies that $\kappa u\in C$ and $\kappa u\in C^{\perp_{L}}$. Since $\pi(\kappa u)=u$ and $(C^{\perp_{L}})_{Res}=(C_{Res})^{\perp}$ by (2) of Theorem 2.1, we have $u\in C_{Res}$ and $u\in(C_{Res})^{\perp}$. Hence, $u\in C_{Res}\cap(C_{Res})^{\perp}$. Therefore, $(C\cap C^{\perp_{L}})_{Res}\subseteq C_{Res}\cap(C_{Res})^{\perp}$. For the converse inclusion, let $v\in C_{Res}\cap(C_{Res})^{\perp}$. This implies that $v\in C_{Res}$ and $v\in(C_{Res})^{\perp}=(C^{\perp_{L}})_{Res}$. Then, by (1) of Theorem 2.1, $\kappa v\in C$ and $\kappa v\in C^{\perp_{L}}$. Hence, $\kappa v\in C\cap C^{\perp_{L}}$. Since $\pi(\kappa v)=v$, we have $v\in(C\cap C^{\perp_{L}})_{Res}$. Therefore, $C_{Res}\cap(C_{Res})^{\perp}\subseteq(C\cap C^{\perp_{L}})_{Res}$. Thus, $(LHull(C))_{Res}=C_{Res}\cap(C_{Res})^{\perp}$.
    On the other hand, let $u\in(LHull(C))_{Tor}$. Then, by (1) of Theorem 2.1, $\zeta u\in C\cap C^{\perp_{L}}$. This implies that $\zeta u\in C$ and $\zeta u\in C^{\perp_{L}}$. Hence, $u\in C_{Tor}$ and $u\in(C^{\perp_{L}})_{Tor}$. Also, by (2) of Theorem 2.1, $u\in C_{Tor}\cap(C_{Res})^{\perp}$. Therefore, $(LHull(C))_{Tor}\subseteq C_{Tor}\cap(C_{Res})^{\perp}$. For the converse inclusion, let $v\in C_{Tor}\cap(C_{Res})^{\perp}$. Then, by (2) of Theorem 2.1, $v\in C_{Tor}$ and $v\in(C^{\perp_{L}})_{Tor}$. Also, by (1) of Theorem 2.1, $\zeta v\in C$ and $\zeta v\in C^{\perp_{L}}$. Consequently, $\zeta v\in C\cap C^{\perp_{L}}$. Hence, $v\in(C\cap C^{\perp_{L}})_{Tor}$. Therefore, $C_{Tor}\cap(C_{Res})^{\perp}\subseteq(C\cap C^{\perp_{L}})_{Tor}$. Thus, $(LHull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}$.
  2. 2.
    The residue code of the right hull can be obtained by the following procedure, analogous to that used for deriving the residue code of the left hull. To derive the torsion code of the right hull, let $u\in(RHull(C))_{Tor}$. Then, by (1) of Theorem 2.1, $\zeta u\in C\cap C^{\perp_{R}}$. This implies that $\zeta u\in C$ and $\zeta u\in C^{\perp_{R}}$. Hence, $u\in C_{Tor}$ and $u\in(C^{\perp_{R}})_{Tor}$. Also, (3) of Theorem 2.1 implies that $u\in C_{Tor}\cap\mathbb{F}_{2}^{n}=C_{Tor}$. Therefore, $(RHull(C))_{Tor}\subseteq C_{Tor}$. For the converse inclusion, let $v\in C_{Tor}$. Since $v\in\mathbb{F}_{2}^{n}$, by (3) of Theorem 2.1, $v\in(C^{\perp_{R}})_{Tor}$. Also, by (1) of Theorem 2.1, $\zeta v\in C$ and $\zeta v\in C^{\perp_{R}}$. Consequently, $\zeta v\in C\cap C^{\perp_{R}}$. Hence, $v\in(C\cap C^{\perp_{R}})_{Tor}$. Therefore, $C_{Tor}\subseteq(C\cap C^{\perp_{R}})_{Tor}$. Thus, $(RHull(C))_{Tor}=C_{Tor}$.
  3. 3.
    This part follows a similar procedure to the above two parts.
∎An important consequence of the above theorem is the following corollary.
Corollary 1
If $C$ is an $E$-linear code, then
$(Hull(C))_{Res}\subseteq Hull(C_{Res})~\text{and}~~(Hull(C))_{Tor}\supseteq Hull(C_{Tor}).$
Furthermore, if $C$ is free, equality holds.
Proof
For an $E$-linear code $C$, by $(3)$ of Theorem 2.3, we have
$(Hull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp}~~\text{and}~~(Hull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}.$
Since $C$ is linear, $C_{Res}\subseteq C_{Tor}$. This implies that $(C_{Tor})^{\perp}\subseteq(C_{Res})^{\perp}$. Hence, $C_{Res}\cap(C_{Tor})^{\perp}\subseteq C_{Res}\cap(C_{Res})^{\perp}$. Therefore, $(Hull(C))_{Res}\subseteq Hull(C_{Res})$. Next, $C_{Tor}\supseteq C_{Res}$ implies that $(C_{Res})^{\perp}\cap C_{Tor}\supseteq(C_{Res})^{\perp}\cap C_{Res}$. Therefore, $(Hull(C))_{Tor}\supseteq Hull(C_{Tor})$. On the other hand, if $C$ is free, then $C_{Res}=C_{Tor}$ implies that
$(Hull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp}=C_{Res}\cap(C_{Res})^{\perp}=Hull(C_{Res}),$
and
$(Hull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}=(C_{Tor})^{\perp}\cap C_{Tor}=Hull(C_{Tor}).$
We know that if $C$ is an $\mathbb{F}_{q}$-linear code, then $(C^{\perp})^{\perp}=C$. Consequently, we have the following results for the three duals of an $E$-linear code.
Theorem 2.4
An $E$-linear code $C$ of length $n$ satisfies the following duality properties:
  1. 1.
    $(C^{\perp_{L}})^{\perp_{L}}=C$ if $C$ is free;
  2. 2.
    $(C^{\perp_{R}})^{\perp_{R}}=C$ if $C_{Res}=\{0\}$ and $C_{Tor}=\mathbb{F}_{2}^{n}$;
  3. 3.
    $(C^{\perp})^{\perp}=C$ always;
  4. 4.
    $(C^{\perp_{L}})^{\perp_{R}}=C$ if $C_{Tor}=\mathbb{F}_{2}^{n}$;
  5. 5.
    $(C^{\perp_{R}})^{\perp_{L}}=C$ if $C$ is free.
Proof
  1. 1.
    If $C$ is an $E$-linear code, then by (2) of Theorem 2.1, we have
    $((C^{\perp_{L}})^{\perp_{L}})_{Res}=((C^{\perp_{L}})_{Res})^{\perp}=((C_{Res})^{\perp})^{\perp}=C_{Res},$
    and
    $((C^{\perp_{L}})^{\perp_{L}})_{Tor}=((C^{\perp_{L}})_{Res})^{\perp}=((C_{Res})^{\perp})^{\perp}=C_{Res}.$
    If $C$ is free, then $C_{Res}=C_{Tor}$ implies that $((C^{\perp_{L}})^{\perp_{L}})_{Res}=C_{Res}$ and $((C^{\perp_{L}})^{\perp_{L}})_{Tor}=C_{Tor}$. Therefore, $(C^{\perp_{L}})^{\perp_{L}}=C$. This completes the proof.
  2. 2.
    If $C$ is an $E$-linear code, then by (3) of Theorem 2.1, we have
    $((C^{\perp_{R}})^{\perp_{R}})_{Res}=((C^{\perp_{R}})_{Tor})^{\perp}=(\mathbb{F}_{2}^{n})^{\perp}=\{0\}~~\text{and}~~((C^{\perp_{R}})^{\perp_{R}})_{Tor}=\mathbb{F}_{2}^{n}.$
    We know that $(C^{\perp_{R}})^{\perp_{R}}=C$ if $((C^{\perp_{R}})^{\perp_{R}})_{Res}=C_{Res}$ and $((C^{\perp_{R}})^{\perp_{R}})_{Tor}=C_{Tor}$. Therefore, $(C^{\perp_{R}})^{\perp_{R}}=C$ if $C_{Res}=\{0\}$ and $C_{Tor}=\mathbb{F}_{2}^{n}$.
  3. 3.
    For an $E$-linear code $C$, (4) of Theorem 2.1 implies that
    $((C^{\perp})^{\perp})_{Res}=((C^{\perp})_{Tor})^{\perp}=((C_{Res})^{\perp})^{\perp}=C_{Res},$
    and
    $((C^{\perp})^{\perp})_{Tor}=((C^{\perp})_{Res})^{\perp}=((C_{Tor})^{\perp})^{\perp}=C_{Tor}.$
    Since $((C^{\perp_{L}})^{\perp_{L}})_{Res}=C_{Res}$ and $((C^{\perp_{L}})^{\perp_{L}})_{Tor}=C_{Tor}$, we have $(C^{\perp})^{\perp}=C$.
  4. 4.
    For an $E$-linear code $C$, (2) and (3) of Theorem 2.1 imply that
    $((C^{\perp_{L}})^{\perp_{R}})_{Res}=((C^{\perp_{L}})_{Tor})^{\perp}=((C_{Res})^{\perp})^{\perp}=C_{Res}~~\text{and}~~((C^{\perp_{L}})^{\perp_{R}})_{Tor}=\mathbb{F}_{2}^{n}.$
    Therefore, $(C^{\perp_{L}})^{\perp_{R}}=C$ if $C_{Tor}=\mathbb{F}_{2}^{n}$.
  5. 5.
    If $C$ is an $E$-linear code, then (2) and (3) of Theorem 2.1 imply that
    $((C^{\perp_{R}})^{\perp_{L}})_{Res}=((C^{\perp_{R}})_{Res})^{\perp}=((C_{Tor})^{\perp})^{\perp}=C_{Tor},$
    and
    $((C^{\perp_{R}})^{\perp_{L}})_{Tor}=((C^{\perp_{R}})_{Res})^{\perp}=((C_{Tor})^{\perp})^{\perp}=C_{Tor}.$
    Therefore, $(C^{\perp_{R}})^{\perp_{L}}=C$ if $C_{Res}=C_{Tor}$. Thus, $(C^{\perp_{R}})^{\perp_{L}}=C$ if $C$ is free.
We know that if $C$ is an $\mathbb{F}_{q}$-linear code, then $Hull(C)=Hull(C^{\perp})$. Correspondingly, we have the following result for three hulls of an $E$-linear code.
Theorem 2.5
An $E$-linear code $C$ of length $n$ satisfies the following properties:
  1. 1.
    $LHull(C)=LHull(C^{\perp_{L}})$ if $C$ is free,
  2. 2.
    $Hull(C)=Hull(C^{\perp})$ always,
  3. 3.
    $RHull(C)=RHull(C^{\perp_{R}})$ if $C_{Res}\cap(C_{Tor})^{\perp}=\{0\}$ and $C_{Tor}=\mathbb{F}_{2}^{n}$.
  4. 4.
    $RHull(C)=LHull(C^{\perp_{R}})$ if $C$ is free,
  5. 5.
    $LHull(C)=RHull(C^{\perp_{L}})$ if $C_{Res}$ is SD and $C$ is free.
Proof
  1. 1.
    If $C$ is a free $E$-linear code, then (1) of Theorem 2.3 implies that
    $(LHull(C^{\perp_{L}}))_{Res}$ $=(C^{\perp_{L}})_{Res}\cap((C^{\perp_{L}})_{Res})^{\perp}$
    $=(C_{Res})^{\perp}\cap((C_{Res})^{\perp})^{\perp}$
    $=(C_{Res})^{\perp}\cap((C_{Res})$
    $=(LHull(C))_{Res},$
    and
    $(LHull(C^{\perp_{L}}))_{Tor}$ $=((C^{\perp_{L}})_{Tor})\cap((C^{\perp_{L}})_{Res})^{\perp}$
    $=C_{Res}\cap(C_{Res})^{\perp}$
    $=C_{Tor}\cap(C_{Res})^{\perp}$
    $=(LHull(C))_{Tor}.$
    Therefore, if $C$ is free, then $LHull(C)=LHull(C^{\perp_{L}})$.
  2. 2.
    From Theorem 2.4 (3), for an $E$-linear code $C$, we have
    $Hull(C^{\perp})=C^{\perp}\cap(C^{\perp})^{\perp}=C^{\perp}\cap C=Hull(C).$
  3. 3.
    Follows from Theorem 2.3(2), for an $E$-linear code $C$, we have
    $(RHull(C^{\perp_{R}}))_{Res}$ $=(C^{\perp_{R}})_{Res}\cap((C^{\perp_{R}})_{Tor})^{\perp}$
    $=(C_{Tor})^{\perp}\cap(\mathbb{F}_{2}^{n})^{\perp}$
    $=(C_{Tor})^{\perp}\cap\{0\}$
    $=\{0\},$
    and
    $(RHull(C^{\perp_{R}}))_{Tor}=(C^{\perp_{R}})_{Tor}=\mathbb{F}_{2}^{n}.$
    Therefore, by (2) of Theorem 2.3, $RHull(C)=RHull(C^{\perp_{R}})$ if $C_{Res}\cap(C_{Tor})^{\perp}=\{0\}$ and $C_{Tor}=\mathbb{F}_{2}^{n}$.
  4. 4.
    If $C$ is a free $E$-linear code, then (1) of Theorem 2.3 implies that
    $(LHull(C^{\perp_{R}}))_{Res}$ $=((C^{\perp_{R}})_{Res})^{\perp}\cap(C^{\perp_{R}})_{Res}$
    $=(C_{Tor})\cap(C_{Tor})^{\perp}$
    $=(C_{Res})\cap(C_{Tor})^{\perp}$
    $=(RHull(C))_{Res},$
    and
    $(LHull(C^{\perp_{R}}))_{Tor}$ $=((C^{\perp_{R}})_{Res})^{\perp}\cap((C^{\perp_{R}})_{Tor})$
    $=((C_{Tor})^{\perp})^{\perp}\cap\mathbb{F}_{2}^{n}$
    $=(C_{Tor})$
    $=(RHull(C))_{Tor}.$
    This shows that $RHull(C)=LHull(C^{\perp_{R}})$.
  5. 5.
    If $C$ is a free $E$-linear code, then (2) of Theorem 2.3 implies that
    $(RHull(C^{\perp_{L}}))_{Res}$ $=((C^{\perp_{L}})_{Tor})^{\perp}\cap(C^{\perp_{L}})_{Res}$
    $=((C_{Res})^{\perp})^{\perp}\cap(C_{Res})^{\perp}$
    $=C_{Res}\cap(C_{Res})^{\perp}$
    $=(LHull(C))_{Res},$
    and
    $(RHull(C^{\perp_{L}}))_{Tor}$ $=((C^{\perp_{L}})_{Tor})$
    $=(C_{Res})^{\perp}$
    $=(C_{Res})\cap(C_{Res})^{\perp}~~~~\hskip 28.45274pt(\because~C_{Res}~~\text{is SD })$
    $=(C_{Res})^{\perp}\cap C_{Tor}~~~~\hskip 36.98866pt(\because~C_{Res}=C_{Tor})$
    $=(LHull(C))_{Tor}.$
    Since $(RHull(C^{\perp_{L}}))_{Res}=(LHull(C))_{Res}$ and $(RHull(C^{\perp_{L}}))_{Tor}=(LHull(C))_{Tor}$, $LHull(C)=RHull(C^{\perp_{L}})$.
Next, we have the following result, which investigates the freeness of the hull codes of a free $E$-linear code.
Theorem 2.6
For a free $E$-linear code $C$, its left hull and two-sided hull codes are also free.
Proof
If $C$ is a free $E$-linear code, then by (1) and (3) of Theorem 2.3, we have
$(LHull(C))_{Res}=C_{Res}\cap(C_{Res})^{\perp},~~(LHull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}$
and
$(Hull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp},~~(Hull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}.$
Since $C$ is free, $C_{Res}=C_{Tor}$. Therefore,
$(LHull(C))_{Res}=(LHull(C))_{Tor}~~\text{and}~~(Hull(C))_{Res}=(Hull(C))_{Tor}.$
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are also free.
Next, we include an example showing that the right hull code of a free $E$-linear code need not be free.
Example 1
Consider the $E$-linear code $C$ with generator matrix
$G=\begin{pmatrix}\kappa&0&0&0\\ 0&\kappa&0&0\end{pmatrix}$.
The residue and torsion codes of the code $C$ have the same generator matrix
$G_{Res}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\end{pmatrix}$.
Therefore, $C$ is free. Next, the dual of the residue code of $C$ is generated by the matrix
$H_{Res}=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\end{pmatrix}$.
Clearly, $C_{Res}\cap(C_{Res})^{\perp}=\{0\}$. Therefore, by (2) of Theorem 2.3, we have
$(RHull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp}=C_{Res}\cap(C_{Res})^{\perp}=\{0\}\neq(RHull(C))_{Tor}.$
Thus, the right hull $RHull(C)$ is not free.
The following result plays a crucial role in our further investigations on hulls.
Theorem 2.7
For a free $E$-linear code $C$, its left hull and two-sided hull codes coincide.
Proof
Since $C$ is free, by (1) and (3) of Theorem 2.3, we have
$(LHull(C))_{Res}=C_{Res}\cap(C_{Res})^{\perp},~~(LHull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}$
and
$(Hull(C))_{Res}=C_{Res}\cap(C_{Tor})^{\perp},~~(Hull(C))_{Tor}=(C_{Res})^{\perp}\cap C_{Tor}.$
Further, the freeness of the $E$-linear code $C$ implies that $C_{Res}=C_{Tor}$. Therefore,
$(LHull(C))_{Res}=(Hull(C))_{Res}~~\text{and}~~(LHull(C))_{Tor}=(Hull(C))_{Tor}.$
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are equal.
According to the above results, the left hull and two-sided hull codes of a free $E$-linear code coincide, without assurance of the freeness of the right hull code. Therefore, from now onward, we focus exclusively on the two-sided hull code of a free $E$-linear code.
For a free $E$-linear code $C$, the following result provides us the generator matrix of its hull code.
Theorem 2.8
If $C$ is a free $E$-linear code with $G$ as a generator matrix of the hull code of $C_{Res}$, then its hull code $Hull(C)$ has a generator matrix $\kappa G$.
Proof
Since $C$ is a free $E$-code, $Hull(C)$ is also free by Theorem 2.6. Next, by Corollary 1, we have $(Hull(C))_{Res}=Hull(C_{Res})$. Moreover, from Theorem $1$ of Kushwaha2 , if $C$ is a free $E$-linear code and $G_{1}$ is a generator matrix of $C_{Res}$, then $\kappa G_{1}$ is a generator matrix of $C$. Therefore, if the hull code of $C_{Res}$ has a generator matrix $G$, then $\kappa G$ is a generator matrix of $Hull(C)$.
Now, we compute the hull-rank of a free $E$-linear code that is needed for the classification of optimal free $E$-linear codes.
Theorem 2.9
If $C$ is a free $E$-linear code, then
$rank(Hull(C))=dim(Hull(C_{Res})).$
Proof
Follows the definition of the rank of a free $E$-linear code and Theorem 2.8.
3 Build-Up Constructions
In this section, we give four build-up construction techniques that construct free $E$-linear codes with a larger length and hull-rank from free $E$-linear codes with a smaller length and hull-rank. Additionally, we give some examples of codes constructed by these build-up construction methods. Note that [$n,k$] represents a free $E$-linear code of length $n$ and rank $k$.
Now, we recall Proposition $1$ of Li18 for our further investigations.
Proposition 1
Let $C$ be an $\mathbb{F}_{q}$-linear $k$-dimensional code with hull-dimension $h$. If $G$ is a generator matrix of $C$ and $G^{T}$ denotes the transpose of $G$, then $h=k-rank(GG^{T})$.
The following result is our first build-up construction method.
Theorem 3.1(Construction I)
Let $C$ be a free $E$-linear [$n,k$]-code. Also, let $G$ (with rows $r_{1},r_{2},\ldots,r_{k}$) and $H$ (with rows $s_{1},s_{2},\ldots,s_{n-k}$) be the generator and parity-check matrices of its residue code $C_{Res}$, respectively. Furthermore, fix a vector $u=(u_{1},u_{2},\ldots,u_{n})\in\mathbb{F}_{2}^{n}$ with $\langle u,u\rangle=1$, and define the following scalars:
$v_{i}=\langle u,r_{i}\rangle~~\text{for}~~1\leq i\leq k,~~\text{and}~~w_{j}=\langle u,s_{j}\rangle~~\text{for}~~1\leq j\leq n-k.$
Now, if the hull-rank of $C$ is $l$, then
  1. (a)
    The $E$-linear code $D$ with generator matrix $G^{\prime}$ given below is a free [$n+2,k+1$]-code with hull-rank $l+1$:
    $G^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&0&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr\kappa v_{1}&\kappa v_{1}&&&\kappa r_{1}&&\\ \kappa v_{2}&\kappa v_{2}&&&\kappa r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ \kappa v_{k}&\kappa v_{k}&&&\kappa r_{k}&&\end{array}\right).$
  2. (b)
    The code $D$ has the following parity-check matrix:
    $H^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&0&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr\kappa w_{1}&\kappa w_{1}&&&\kappa s_{1}&&\\ \kappa w_{2}&\kappa w_{2}&&&\kappa s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ \kappa w_{n-k}&\kappa w_{n-k}&&&\kappa s_{n-k}&&\end{array}\right).$
Proof
  1. (a)
    By the form of $G^{\prime}$, it is clear that it generates a free $E$-linear code, and its residue code has the generator matrix
    $G_{1}=\left(\begin{array}[]{cc|ccccc}1&0&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr v_{1}&v_{1}&&&r_{1}&&\\ v_{2}&v_{2}&&&r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ v_{k}&v_{k}&&&r_{k}&&\end{array}\right).$
    Then, we have
    $G_{1}G_{1}^{T}=\left(\begin{array}[]{ccccccc}0&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&GG^{T}&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right).$
    Therefore, $rank(G_{1}G_{1}^{T})=rank(GG^{T})$. Since $C$ is free, by Theorem 2.9, we have $dim(Hull(C_{Res}))=rank(Hull(C))=l$. Also, by Proposition 1, $rank(GG^{T})=k-l$ and hence $rank(G_{1}G_{1}^{T})=k-l$. Again, Proposition 1 implies that the hull-dimension of $D_{Res}$ is given by $(k+1)-rank(G_{1}G_{1}^{T})=(k+1)-(k-l)=l+1$. Since, for a free $E$-linear code $D$, $rank(D)=dim(D_{Res})$, we have $rank(Hull(D))=dim(Hull(D_{Res}))=l+1$. This completes the proof of the first part.
  2. (b)
    We have
    $\pi(H^{\prime})=H_{1}=\left(\begin{array}[]{cc|ccccc}1&0&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr w_{1}&w_{1}&&&s_{1}&&\\ w_{2}&w_{2}&&&s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ w_{n-k}&w_{n-k}&&&s_{n-k}&&\end{array}\right).$
    Then, we have
    $G_{1}H_{1}^{T}=\left(\begin{array}[]{ccccccc}0&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&GH^{T}&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right)=\boldsymbol{0}.$
    The dimension of the code generated by $H_{1}$ is (nk)+1(n-k)+1, since the first row of $H_{1}$ is not a linear combination of the other rows of $H_{1}$. Further,
    $dim\langle H_{1}\rangle=(n-k)+1=(n+2)-(k+1)=dim((D^{\perp})_{Res}).$
    Therefore, $H_{1}$ is a parity-check matrix of $(D)_{Res}$. Thus, $H^{\prime}$ is a parity-check matrix of $D$.
Next, we give an example which utilizes a free $E$-linear [$6,4$]-code with hull-rank $2$ to construct a free [$8,5$]-code over $E$ with hull-rank $3$ by Construction I.
Example 2
Consider the free $E$-linear code $C$ with generator matrix
$N=\left(\begin{array}[]{cccccc}\kappa&0&0&0&\kappa&0\\ 0&\kappa&0&0&\kappa&\kappa\\ 0&0&\kappa&0&0&\kappa\\ 0&0&0&\kappa&\kappa&\kappa\end{array}\right).$
Clearly, $C$ is a [$6,4$]-code with hull-rank $2$, and $C_{Res}$ has a generator matrix
$G=\left(\begin{array}[]{cccccc}1&0&0&0&1&0\\ 0&1&0&0&1&1\\ 0&0&1&0&0&1\\ 0&0&0&1&1&1\end{array}\right).$
Next, if we take $u=(1,0,0,1,0,1)$, then $\langle u,u\rangle=1$ and $v_{1}=1,v_{2}=1,v_{3}=1$ and $v_{4}=0$. Now, consider the code $D$ with the following generator matrix
$G^{\prime}=\left(\begin{array}[]{cc|cccccc}\kappa&0&\kappa&0&0&\kappa&0&\kappa\\ \hline\cr\kappa&\kappa&&&&&&\\ \kappa&\kappa&&&N&&&\\ \kappa&\kappa&&&&&&\\ 0&0&&&&&&\end{array}\right).$
Then, by Construction I, $D$ is a free [$8,5$]-code over $E$ with hull-rank $3$.
Now, we have our second build-up construction method as follows.
Theorem 3.2(Construction II)
Let $C$ be a free $E$-linear [$n,k$]-code. Let $G$ (with rows $r_{1},r_{2},\ldots,r_{k}$) and $H$ (with rows $s_{1},s_{2},\ldots,s_{n-k}$) be the generator and parity-check matrices of its residue code $C_{Res}$, respectively. Furthermore, fix a vector $u=(u_{1},u_{2},\ldots,u_{n})\in\mathbb{F}_{2}^{n}$ with $\langle u,u\rangle=0$ and $\langle u,r_{i}\rangle=0~~\text{for}~~1\leq i\leq k$. Also, define $w_{j}=\langle u,s_{j}\rangle~~\text{for}~~1\leq j\leq n-k$. With these assumptions, if the hull-rank of $C$ is $l$, then
  1. (a)
    The code $D$ with the generator matrix
    $G^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&\kappa&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr 0&0&&&\kappa r_{1}&&\\ 0&0&&&\kappa r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ 0&0&&&\kappa r_{k}&&\end{array}\right)$
    is a free $E$-linear [$n+2,k+1$]-code with hull-rank $l+1$.
  2. (b)
    The code $D$ has the following parity-check matrix:
    $H^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&\kappa&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr 0&\kappa w_{1}&&&\kappa s_{1}&&\\ 0&\kappa w_{2}&&&\kappa s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ 0&\kappa w_{n-k}&&&\kappa s_{n-k}&&\end{array}\right).$
Proof
  1. (a)
    It is clear that $G^{\prime}$ generates a free $E$-linear code, and its residue code has the generator matrix
    $G_{1}=\left(\begin{array}[]{cc|ccccc}1&1&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr 0&0&&&r_{1}&&\\ 0&0&&&r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ 0&0&&&r_{k}&&\end{array}\right)$
    Then, we have
    $G_{1}G_{1}^{T}=\left(\begin{array}[]{ccccccc}0&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&GG^{T}&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right).$
    The proof follows similar arguments to part $(a)$ of Construction I.
  2. (b)
    We have
    $\pi(H^{\prime})=H_{1}=\left(\begin{array}[]{cc|ccccc}1&1&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr 0&w_{1}&&&s_{1}&&\\ 0&w_{2}&&&s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ 0&w_{n-k}&&&s_{n-k}&&\end{array}\right).$
    Then
    $G_{1}H_{1}^{T}=\left(\begin{array}[]{ccccccc}0&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&GH^{T}&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right)=\boldsymbol{0}.$
    The remainder of the proof is parallel to the part $(b)$ of Construction I.
The following example utilizes a free $E$-linear [$9,4$]-code with hull-rank $4$ to construct a free [$11,5$]-code over $E$ with hull-rank $5$ by Construction II.
Example 3
Let $C$ be an $E$-linear code with generator matrix
$N=\left(\begin{array}[]{ccccccccc}\kappa&0&0&0&\kappa&\kappa&0&\kappa&0\\ 0&\kappa&0&0&0&\kappa&0&\kappa&\kappa\\ 0&0&\kappa&0&\kappa&\kappa&0&0&\kappa\\ 0&0&0&\kappa&\kappa&0&0&\kappa&\kappa\end{array}\right).$
Then, $C_{Res}$ is generated by the matrix
$G=\left(\begin{array}[]{ccccccccc}1&0&0&0&1&1&0&1&0\\ 0&1&0&0&0&1&0&1&1\\ 0&0&1&0&1&1&0&0&1\\ 0&0&0&1&1&0&0&1&1\end{array}\right).$
Clearly, $C$ is a free [$9,4$]-code over $E$ with hull-rank $4$. If we take $u=(1,0,0,0,1,1,0,1,0)$, then $\langle u,u\rangle=0$ and $\langle u,r_{i}\rangle=0$ for $1\leq i\leq 4$. Now, let $D$ be the $E$-linear code generated by the following matrix
$G^{\prime}=\left(\begin{array}[]{cc|ccccccccc}\kappa&\kappa&\kappa&0&0&0&\kappa&\kappa&0&\kappa&0\\ \hline\cr 0&0&&&&&&&&&\\ 0&0&&&&N&&&&&\\ 0&0&&&&&&&&&\\ 0&0&&&&&&&&&\end{array}\right).$
Then, by Construction II, $D$ is a free [$11,5$]-code over $E$ with hull-rank $5$.
The following result is our third build-up construction method.
Theorem 3.3(Construction III)
Let $C$ be a free $E$-linear [$n,k$]-code. Also, let $G$ (with rows $r_{1},r_{2},\ldots,r_{k}$) and $H$ (with rows $s_{1},s_{2},\ldots,s_{n-k}$) be the generator and parity-check matrices of its residue code $C_{Res}$, respectively. Furthermore, fix a vector $u=(u_{1},u_{2},\ldots,u_{n})\in\mathbb{F}_{2}^{n}$ with $\langle u,u\rangle=0$, and define the following scalars:
$v_{i}=\langle u,r_{i}\rangle~~\text{for}~~1\leq i\leq k,~~\text{and}~~w_{j}=\langle u,s_{j}\rangle~~\text{for}~~1\leq j\leq n-k.$
Now, assume that not all $v_{i}$’s are zero and hull-rank of $C$ is $l$, then
  1. (a)
    The $E$-linear code $D$ with generator matrix $G^{\prime}$ given below is a free [$n+2,k+1$]-code with hull-rank $l,l+1$ or $l+2$:
    $G^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&\kappa&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr\kappa v_{1}&0&&&\kappa r_{1}&&\\ \kappa v_{1}&0&&&\kappa r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ \kappa v_{1}&0&&&\kappa r_{k}&&\end{array}\right).$
  2. (b)
    The code $D$ has the following parity-check matrix:
    $H^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&\kappa&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr 0&\kappa w_{1}&&&\kappa s_{1}&&\\ 0&\kappa w_{2}&&&\kappa s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ 0&\kappa w_{n-k}&&&\kappa s_{n-k}&&\end{array}\right).$
Proof
  1. (a)
    Follows the form of $G^{\prime}$, it generates a free $E$-linear code, and its residue code has the generator matrix
    $G_{1}=\left(\begin{array}[]{cc|ccccc}1&1&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr v_{1}&0&&&r_{1}&&\\ v_{1}&0&&&r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ v_{1}&0&&&r_{k}&&\end{array}\right).$
    Then, applying elementary row operations on $G_{1}$, we get
    $G^{\prime}_{1}=\left(\begin{array}[]{cc|ccccc}1&1&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr 1&0&&&&&\\ 0&0&&&G_{2}&&\\ \vdots&\vdots&&&&&\\ 0&0&&&&&\end{array}\right),$
    where $\langle G_{2}\rangle=\langle G\rangle$. Now, we have
    $G^{\prime}_{1}(G^{\prime}_{1})^{T}=\left(\begin{array}[]{ccccccc}0&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&M&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right),$
    where $M=(1,0,\ldots,0)^{T}(1,0,\ldots,0)+G_{2}G_{2}^{T}$. We know that if $M_{1}$ and $M_{2}$ are two matrices, then $rank(M_{1}+M_{2})\leq rank(M_{1})+rank(M_{2})$. Therefore,
    $rank(M)$ $\leq rank((1,0,\ldots,0)^{T}(1,0,\ldots,0))+rank(G_{2}G_{2}^{T})$
    $\leq 1+rank(G_{2}G_{2}^{T}).$
    We observe that $(1,0,\ldots,0)^{T}(1,0,\ldots,0)$ affects exclusively the first row of $G_{2}G_{2}^{T}$. Hence, $rank(M)$ can be $rank(G_{2}G_{2}^{T})$, $rank(G_{2}G_{2}^{T})-1$ or $rank(G_{2}G_{2}^{T})+1$. Then, the following cases arise:
    Case 1: Let $rank(M)=rank(G_{2}G_{2}^{T})$. Then
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)$ $=rank(G_{2}G_{2}^{T})=rank(GG^{T})=k-l.$
    Therefore, hull-dimension of $D_{Res}$ is (k+1)(kl)=l+1(k+1)-(k-l)=l+1. Thus, by Theorem 2.9, the hull-rank of the code $D$ is $l+1$ in this case.
    Case 2: Let $rank(M)=rank(G_{2}G_{2}^{T})-1$. Then
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)=rank(G_{2}G_{2}^{T})-1=rank(GG^{T})-1=k-l-1.$
    Therefore, hull-dimension of $D_{Res}$ is (k+1)(kl1)=l+2(k+1)-(k-l-1)=l+2. Thus, by Theorem 2.9, the hull-rank of the code $D$ is $l+2$ in this case.
    Case 3: Let $rank(M)=rank(G_{2}G_{2}^{T})+1$. Then
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)=rank(G_{2}G_{2}^{T})+1=rank(GG^{T})+1=k-l+1.$
    Therefore, $D_{Res}$ has hull-dimension (k+1)(kl+1)=l(k+1)-(k-l+1)=l. Hence, by Theorem 2.9, the hull-rank of the code $D$ is $l$ in this case.
  2. (b)
    This part follows similar approach as the part $(b)$ of Construction I.
Now, we have a supporting example for a free $E$-linear [$10,6$]-code with hull-rank $1$ to construct a free [$12,7$]-code over $E$ with hull-rank $3$ by Construction III.
Example 4
Let $C$ be a free $E$-linear [$10,6$]-code with hull-rank $1$, and $C_{Res}$ has a generator matrix
$G=\left(\begin{array}[]{cccccccccc}1&0&0&0&0&0&1&0&0&1\\ 0&1&0&0&0&0&0&1&0&1\\ 0&0&1&0&0&0&1&1&0&1\\ 0&0&0&1&0&0&1&1&1&1\\ 0&0&0&0&1&0&1&1&0&0\\ 0&0&0&0&0&1&0&0&1&1\end{array}\right).$
If $u=(1,1,1,1,1,1,1,1,1,1)$, then $\langle u,u\rangle=0$, and $v_{1}=v_{2}=v_{4}=v_{5}=v_{6}=1,v_{3}=0$. Now, consider the code $D$ with generator matrix
$G^{\prime}=\left(\begin{array}[]{cc|cccccccccc}\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa\\ \hline\cr\kappa&0&&&&&&&&&&\\ \kappa&0&&&&&&&&&&\\ 0&0&&&&\kappa G&&&&&&\\ \kappa&0&&&&&&&&&&\\ \kappa&0&&&&&&&&&&\\ \kappa&0&&&&&&&&&&\end{array}\right).$
Then, by Construction III, $D$ is a free [$12,7$]-code over $E$ with hull-rank $3$.
The following result is our fourth build-up construction method.
Theorem 3.4(Construction IV)
Let $C$ be a free $E$-linear [$n,k$]-code. Also, let $G$ (with rows $r_{1},r_{2},\ldots,r_{k}$) and $H$ (with rows $s_{1},s_{2},\ldots,s_{n-k}$) be the generator and parity-check matrices of its residue code $C_{Res}$, respectively. Furthermore, fix a vector $u=(u_{1},u_{2},\ldots,u_{n})\in\mathbb{F}_{2}^{n}$ with $\langle u,u\rangle=0$, and define the following scalars:
$v_{i}=\langle u,r_{i}\rangle~~\text{for}~~1\leq i\leq k,~~\text{and}~~w_{j}=\langle u,s_{j}\rangle~~\text{for}~~1\leq j\leq n-k.$
Assume the hull-rank of $C$ is $l$, then
  1. (a)
    The code $D$ with the generator matrix
    $G^{\prime}=\left(\begin{array}[]{cc|ccccc}\kappa&0&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr\kappa v_{1}&\kappa v_{1}&&&\kappa r_{1}&&\\ \kappa v_{2}&\kappa v_{2}&&&\kappa r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ \kappa v_{k}&\kappa v_{k}&&&\kappa r_{k}&&\end{array}\right)$
    is a free $E$-linear [$n+2,k+1$]-code with hull-rank $l$.
  2. (b)
    The code D has the following parity-check matrix:
    $H^{\prime}=\left(\begin{array}[]{cc|ccccc}0&\kappa&\kappa u_{1}&\kappa u_{2}&\cdots\cdots&&\kappa u_{n}\\ \hline\cr\kappa w_{1}&\kappa w_{1}&&&\kappa s_{1}&&\\ \kappa w_{2}&\kappa w_{2}&&&\kappa s_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ \kappa w_{n-k}&\kappa w_{n-k}&&&\kappa s_{n-k}&&\end{array}\right).$
Proof
  1. (a)
    It is clear that $G^{\prime}$ generates a free $E$-linear code, and its residue code has the generator matrix
    $G_{1}=\left(\begin{array}[]{cc|ccccc}1&0&u_{1}&u_{2}&\cdots\cdots&&u_{n}\\ \hline\cr v_{1}&v_{1}&&&r_{1}&&\\ v_{2}&v_{2}&&&r_{2}&&\\ \vdots&\vdots&&&\vdots&&\\ v_{k}&v_{k}&&&r_{k}&&\end{array}\right).$
    Then
    $G_{1}G_{1}^{T}=\left(\begin{array}[]{ccccccc}1&0&0&0&\cdots\cdots&&0\\ 0&&&&&&\\ 0&&&GG^{T}&&&\\ \vdots&&&&&&\\ 0&&&&&&\end{array}\right).$
    Hence, $rank(G_{1}G_{1}^{T})=rank(GG^{T})+1=k-l+1$. Again, Proposition 1 implies that the hull-dimension of $D_{Res}$ is given by $(k+1)-rank(G_{1}G_{1}^{T})=(k+1)-(k-l+1)=l$. Therefore, $rank(Hull(D))=dim(Hull(D_{Res}))=l$.
  2. (b)
    This part can be proved similar to the part $(b)$ of previous constructions.
This section concludes with the following example, which illustrates Construction IV.
Example 5
Let $C$ be a free [$10,5$]-code over $E$ with hull-rank $5$, and $C_{Res}$ has a generator matrix
$G=\left(\begin{array}[]{cccccccccc}1&0&0&0&0&0&0&1&0&0\\ 0&1&0&0&0&0&1&0&0&0\\ 0&0&1&0&0&0&0&0&0&1\\ 0&0&0&1&0&0&0&0&1&0\\ 0&0&0&0&1&1&0&0&0&0\end{array}\right).$
If we take $u=(1,0,1,1,0,1,0,0,0,0)$, then $\langle u,u\rangle=0$, and $v_{1}=v_{3}=v_{4}=v_{5}=v_{6}=1,v_{2}=0$. Now, consider the code $D$ with generator matrix
$G^{\prime}=\left(\begin{array}[]{cc|cccccccccc}\kappa&0&\kappa&0&\kappa&\kappa&0&\kappa&0&0&0&0\\ \hline\cr\kappa&\kappa&&&&&&&&&&\\ 0&0&&&&&&&&&&\\ \kappa&\kappa&&&&&\kappa G&&&&&\\ \kappa&\kappa&&&&&&&&&&\\ \kappa&\kappa&&&&&&&&&&\\ \end{array}\right).$
Then, by Construction IV, $D$ is a free [$12,6$]-code over $E$ with hull-rank $5$.
4 Classification
This section investigates the permutation equivalence of free $E$-linear codes and classifies the optimal free $E$-linear codes with a fixed hull-rank. We also study the hull-variation problem for the free $E$-linear codes.
Definition 7(Permutation-equivalent codes)
A code $C$ is called permutation-equivalent to a code $D$ if $C$ can be obtained from $D$ by a suitable coordinate permutation.
The following result from Alah23 investigates the permutation equivalence of two free $E$-linear codes.
Theorem 4.1
( Alah23 , Theorem $16$) Two free $E$-linear codes $C$ and $D$ are permutation-equivalent if and only if $C_{Res}$ and $D_{Res}$ are permutation-equivalent.
To study the hull-variation problem for the free $E$-linear codes, we have the following.
Theorem 4.2
Let $C$ be a free $E$-linear [$n,k$]-code and $\alpha$ be a permutation. Then
  1. 1.
    $\langle\alpha(w),\alpha(z)\rangle=\langle w,z\rangle$ for all $w,z\in E^{n}$,
  2. 2.
    $\alpha(C^{\perp})=\alpha(C)^{\perp}$.
Proof
  1. 1.
    Let $w=(w_{1},w_{2},\ldots,w_{n}),z=(z_{1},z_{2},\ldots,z_{n})\in E^{n}$ and $\alpha$ be a permutation. Then
    $\alpha(w)=(w_{\alpha(1)},w_{\alpha(2)},\ldots,w_{\alpha(n)})~~\text{and}~~~\alpha(z)=(z_{\alpha(1)},z_{\alpha(2)},\ldots,z_{\alpha(n)}).$
    Therefore,
    $\langle\alpha(w),\alpha(z)\rangle=\sum_{i=1}^{n}w_{\alpha(i)}z_{\alpha(i)}=\sum_{i=1}^{n}w_{i}z_{i}=\langle w,z\rangle.$
  2. 2.
    Let $w\in C^{\perp}$ and $\alpha$ be a permutation. Then $\alpha(w)\in\alpha(C^{\perp})$. Now, for any $z\in\alpha(C)$, there exists a codeword $x\in C$ such that $\alpha(x)=z$. Hence, by the first part, we have
    $\langle z,\alpha(w)\rangle=\langle\alpha(x),\alpha(w)\rangle=\langle x,w\rangle=0.$
    Therefore, $\alpha(w)\in\alpha(C)^{\perp}$, and hence $\alpha(C^{\perp})\subseteq(\alpha(C))^{\perp}$. Further, for the free [$n,k$]-code $C$ over $E$,
    $rank(\alpha(C^{\perp}))=rank(C^{\perp})=n-rank(C)=n-rank(\alpha(C))=rank(\alpha(C)^{\perp}).$
    This implies that $\alpha(C^{\perp})=\alpha(C)^{\perp}$.
We now investigate the hull-variation problem for the free $E$-linear codes.
Theorem 4.3
The hull-ranks of any two permutation-equivalent free $E$-linear codes are identical.
Proof
Let $C$ be a free $E$-linear code and $\alpha$ be an arbitrary permutation. Then, Theorem 4.2 implies that
$\alpha(Hull(C))=\alpha(C\cap C^{\perp})=\alpha(C)\cap\alpha(C^{\perp})=\alpha(C)\cap\alpha(C)^{\perp}=Hull(\alpha(C)).$
Therefore,
$rank(Hull(\alpha(C)))=rank(\alpha(Hull(C)))=rank(Hull(C)).$
Definition 8($l_{i}$-optimal free code)
A free $E$-linear [$n,k$]-code $C$ with hull-rank $l=i$ is called an $l_{i}$-optimal code if the minimum distance of the code $C$ is maximum among all the free [$n,k$]-codes over $E$ with hull-rank $i$.
Here, we classify optimal free $E$-linear codes with a fixed hull-rank. Since LCD $E$-codes are free $E$-linear codes whose hull has rank 0, and their classification is already provided in Kushwaha2 . Therefore, we classify only the free $E$-linear codes with non-zero hull-ranks. We use Theorem $1$ of Kushwaha2 to get all the free $E$-linear codes, and then use Theorem 4.1 to find all the permutation-inequivalent free linear codes over $E$. Further, Theorem 2.9 helps to find their hull-ranks. Later, we use the Definition 8 to determine their optimality and list them in Tables 3, 3, 5, 5, 7, 7, 9, 9 and 10. In the tables, $I_{m}$ is the identity matrix of order $m$. Here, all computations have been carried out using MAGMA software Magma .
Table 2: $l_{1}$-optimal free codes (Part I).
Generator Matrix $[n,k,d]$
$\begin{pmatrix}\kappa&\kappa\end{pmatrix}$ $[2,1,2]$
$\begin{pmatrix}\kappa&\kappa&0\end{pmatrix}$ $[3,1,2]$
$\begin{pmatrix}\kappa&0&\kappa\\ 0&\kappa&0\end{pmatrix}$ $[3,2,1]$
$\begin{pmatrix}\kappa&\kappa&\kappa\end{pmatrix}$ $[4,1,4]$
$\begin{pmatrix}\kappa&0&0&0\\ 0&\kappa&\kappa&0\end{pmatrix}$ $[4,2,1]$
$\begin{pmatrix}&\kappa\\ \kappa I_{3}&\kappa\\ &\kappa\end{pmatrix}$ $[4,3,2]$
$\begin{pmatrix}\kappa&\kappa&\kappa&0&\kappa\end{pmatrix}$ $[5,1,4]$
$\begin{pmatrix}\kappa&0&0&\kappa&\kappa\\ 0&\kappa&\kappa&\kappa&0\end{pmatrix}$ $[5,2,3]$
$\begin{pmatrix}&0&\kappa\\ \kappa I_{3}&\kappa&\kappa\\ &\kappa&\kappa\end{pmatrix}$ $[5,3,2]$
$\begin{pmatrix}&0&\kappa\\ \kappa I_{3}&\kappa&0\\ &0&\kappa\end{pmatrix}$ $[5,3,2]$
$\begin{pmatrix}&0&\kappa\\ \kappa I_{3}&0&\kappa\\ &0&\kappa\end{pmatrix}$ $[5,3,2]$
$\begin{pmatrix}&\kappa\\ \kappa I_{4}&0\\ &0\\ &0\end{pmatrix}$ $[5,4,1]$
$\begin{pmatrix}&\kappa\\ \kappa I_{4}&\kappa\\ &\kappa\\ &0\end{pmatrix}$ $[5,4,1]$
$\begin{pmatrix}\kappa&\kappa&\kappa&\kappa&\kappa&\kappa\end{pmatrix}$ $[6,1,6]$
$\begin{pmatrix}\kappa&0&\kappa&0&0&\kappa\\ 0&\kappa&\kappa&\kappa&0&0\end{pmatrix}$ $[6,2,3]$
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{3}&0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&\kappa&0&\kappa\\ \kappa I_{3}&0&0&\kappa\\ &0&\kappa&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&0&\kappa&\kappa\\ \kappa I_{3}&0&0&\kappa\\ &0&\kappa&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&\kappa&0&0\\ \kappa I_{3}&0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&\kappa&\kappa&\kappa\\ \kappa I_{3}&0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&\kappa&0&0\\ \kappa I_{3}&0&0&\kappa\\ &0&\kappa&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&0&\kappa&0\\ \kappa I_{3}&0&0&\kappa\\ &\kappa&\kappa&\kappa\end{pmatrix}$ $[6,3,2]$
$\begin{pmatrix}&0&\kappa\\ \kappa I_{4}&\kappa&0\\ &\kappa&\kappa\\ &0&0\end{pmatrix}$ [$6,4,1$]
Table 3: $l_{1}$-optimal free codes (Part II).
Generator Matrix [$n,k,d$]
$\begin{pmatrix}&0&\kappa\\ \kappa I_{4}&0&\kappa\\ &0&0\\ &0&\kappa\end{pmatrix}$ [$6,4,1$]
$\begin{pmatrix}&0&\kappa\\ \kappa I_{4}&0&0\\ &0&0\\ &0&0\end{pmatrix}$ [$6,4,1$]
$\begin{pmatrix}&0&\kappa\\ \kappa I_{4}&\kappa&0\\ &\kappa&0\\ &0&0\end{pmatrix}$ [$6,4,1$]
$\begin{pmatrix}&\kappa\\ &\kappa\\ \kappa I_{5}&\kappa\\ &\kappa\\ &\kappa\end{pmatrix}$ [$6,5,2$]
$\begin{pmatrix}\kappa&\kappa&\kappa&\kappa&0&\kappa&\kappa\end{pmatrix}$ [$7,1,6$]
$\begin{pmatrix}\kappa&0&\kappa&\kappa&0&\kappa&\kappa\\ 0&\kappa&0&\kappa&\kappa&\kappa&\kappa\end{pmatrix}$ [$7,2,4$]
$\begin{pmatrix}&\kappa&\kappa&0&\kappa\\ \kappa I_{3}&\kappa&\kappa&0&0\\ &\kappa&0&\kappa&\kappa\end{pmatrix}$ [$7,3,3$]
$\begin{pmatrix}&0&\kappa&0\\ \kappa I_{4}&\kappa&0&\kappa\\ &0&\kappa&0\\ &0&\kappa&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&\kappa&0\\ \kappa I_{4}&\kappa&0&\kappa\\ &0&\kappa&0\\ &0&\kappa&\kappa\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&\kappa&\kappa&0\\ \kappa I_{4}&\kappa&0&\kappa\\ &0&\kappa&0\\ &0&\kappa&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&\kappa&\kappa&0\\ \kappa I_{4}&\kappa&0&\kappa\\ &\kappa&0&0\\ &\kappa&0&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&\kappa&\kappa\\ \kappa I_{4}&\kappa&\kappa&\kappa\\ &0&\kappa&0\\ &\kappa&\kappa&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&\kappa&0\\ \kappa I_{4}&\kappa&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&\kappa\\ &0&\kappa\\ \kappa I_{5}&0&\kappa\\ &0&\kappa\\ &0&\kappa\\ \end{pmatrix}$ [$7,5,2$]
$\begin{pmatrix}&\kappa&0\\ &0&\kappa\\ \kappa I_{5}&\kappa&0\\ &\kappa&0\\ &0&\kappa\\ \end{pmatrix}$ [$7,5,2$]
$\begin{pmatrix}&\kappa&0\\ &0&\kappa\\ \kappa I_{5}&0&\kappa\\ &0&\kappa\\ &0&\kappa\\ \end{pmatrix}$ [$7,5,2$]
$\begin{pmatrix}&\kappa&\kappa\\ &0&\kappa\\ \kappa I_{5}&\kappa&\kappa\\ &0&\kappa\\ &0&\kappa\\ \end{pmatrix}$ [$7,5,2$]
Table 4: $l_{1}$-optimal free codes (Part III).
Generator Matrix $[n,k,d]$
$\begin{pmatrix}&\kappa&\kappa\\ &\kappa&0\\ \kappa I_{5}&\kappa&\kappa\\ &0&\kappa\\ &0&\kappa\\ \end{pmatrix}$ [$7,5,2$]
$\begin{pmatrix}&0\\ &0\\ \kappa I_{6}&0\\ &0\\ &0\\ &\kappa\end{pmatrix}$ [$7,6,1$]
$\begin{pmatrix}&\kappa\\ &\kappa\\ \kappa I_{6}&\kappa\\ &\kappa\\ &\kappa\\ &0\end{pmatrix}$ [$7,6,1$]
$\begin{pmatrix}&\kappa\\ &0\\ \kappa I_{6}&\kappa\\ &0\\ &\kappa\\ &0\end{pmatrix}$ [$7,6,1$]
$\begin{pmatrix}\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\kappa&\end{pmatrix}$ [$8,1,8$]
$\begin{pmatrix}\kappa&0&\kappa&0&\kappa&\kappa&0&0\\ 0&\kappa&\kappa&\kappa&0&\kappa&0&\kappa\end{pmatrix}$ [$8,2,4$]
$\begin{pmatrix}&\kappa&\kappa&\kappa&0&0\\ \kappa I_{3}&\kappa&\kappa&0&\kappa&0\\ &0&\kappa&\kappa&0&\kappa\end{pmatrix}$ [$8,3,4$]
$\begin{pmatrix}&\kappa&0&0&\kappa\\ \kappa I_{4}&\kappa&\kappa&0&0\\ &\kappa&0&0&\kappa\\ &\kappa&0&\kappa&0\end{pmatrix}$ [$8,4,3$]
$\begin{pmatrix}&\kappa&0&\kappa\\ &0&\kappa&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&0&\kappa\\ &0&\kappa&0\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &\kappa&0&\kappa\\ \kappa I_{5}&0&\kappa&\kappa\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&0&0\\ &0&0&\kappa\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&0&0\\ &\kappa&\kappa&\kappa\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&0&\kappa\\ &0&\kappa&0\\ \kappa I_{5}&0&0&\kappa\\ &\kappa&\kappa&\kappa\\ &\kappa&\kappa&0\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &0&\kappa&0\\ \kappa I_{5}&0&\kappa&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&0&0\\ &0&\kappa&\kappa\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
Table 5: $l_{1}$-optimal free codes (Part IV).
Generator Matrix [$n,k,d$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &0&0&\kappa\\ \kappa I_{5}&0&\kappa&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&\kappa&0\\ &\kappa&0&\kappa\\ \kappa I_{5}&\kappa&0&0\\ &\kappa&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&\kappa&0\\ &0&\kappa&\kappa\\ \kappa I_{5}&\kappa&0&0\\ &\kappa&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &0&0&\kappa\\ \kappa I_{5}&\kappa&0&0\\ &\kappa&0&0\\ &0&\kappa&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&\kappa&\kappa\\ &0&\kappa&0\\ \kappa I_{5}&\kappa&0&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &\kappa&0&\kappa\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &0&0&\kappa\\ \kappa I_{5}&0&\kappa&0\\ &0&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&0&\kappa\\ &0&0&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&0&\kappa\\ &\kappa&\kappa&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &\kappa&0&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &0&0&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &\kappa&\kappa&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&\kappa\\ \kappa I_{6}&0&0\\ &0&0\\ &0&0\\ &\kappa&\kappa\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&0&\kappa\\ &0&0\\ \kappa I_{6}&0&0\\ &0&0\\ &0&0\\ &0&0\end{pmatrix}$ [$8,6,1$]
Table 6: $l_{1}$-optimal free codes (Part V).
Generator Matrix $[n,k,d]$
$\begin{pmatrix}&0&\kappa\\ &0&\kappa\\ \kappa I_{6}&0&\kappa\\ &0&\kappa\\ &0&\kappa\\ &0&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&\kappa&0\\ &0&\kappa\\ \kappa I_{6}&0&\kappa\\ &0&\kappa\\ &0&0\\ &0&\kappa\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&\kappa\\ \kappa I_{6}&0&0\\ &0&\kappa\\ &0&\kappa\\ &\kappa&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&0&\kappa\\ &0&\kappa\\ \kappa I_{6}&0&0\\ &0&\kappa\\ &0&0\\ &0&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&\kappa&\kappa\\ &0&\kappa\\ \kappa I_{6}&\kappa&0\\ &0&\kappa\\ &\kappa&\kappa\\ &0&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&0\\ \kappa I_{6}&0&\kappa\\ &0&\kappa\\ &0&0\\ &\kappa&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&\kappa&0\\ &0&\kappa\\ \kappa I_{6}&0&\kappa\\ &0&0\\ &0&\kappa\\ &0&0\end{pmatrix}$ [$8,6,1$]
$\begin{pmatrix}&\kappa\\ &\kappa\\ &\kappa\\ \kappa I_{6}&\kappa\\ &\kappa\\ &\kappa\\ &\kappa\\ \end{pmatrix}$ [$8,7,2$]
Table 7: $l_{2}$-optimal free codes (Part I).
Generator Matrix $[n,k,d]$
$\begin{pmatrix}\kappa&0&\kappa&0\\ 0&\kappa&0&\kappa\end{pmatrix}$ [$4,2,2$]
$\begin{pmatrix}\kappa&0&0&\kappa&0\\ 0&\kappa&\kappa&0&0\end{pmatrix}$ [$5,2,2$]
$\begin{pmatrix}&0&\kappa\\ \kappa I_{3}&\kappa&0\\ &0&0\end{pmatrix}$ [$5,3,1$]
$\begin{pmatrix}\kappa&0&\kappa&0&\kappa&\kappa\\ 0&\kappa&\kappa&\kappa&0&\kappa\end{pmatrix}$ [$6,2,4$]
$\begin{pmatrix}&\kappa&\kappa&\kappa\\ \kappa I_{3}&\kappa&0&\kappa\\ &0&\kappa&\kappa\end{pmatrix}$ [$6,3,3$]
$\begin{pmatrix}&0&\kappa\\ \kappa I_{4}&\kappa&0\\ &\kappa&0\\ &\kappa&0\end{pmatrix}$ [$6,4,2$]
$\begin{pmatrix}&\kappa&0\\ \kappa I_{4}&\kappa&\kappa\\ &0&\kappa\\ &\kappa&\kappa\end{pmatrix}$ [$6,4,2$]
$\begin{pmatrix}\kappa&0&0&\kappa&\kappa&0&\kappa\\ 0&\kappa&\kappa&\kappa&\kappa&0&0\end{pmatrix}$ [$7,2,4$]
$\begin{pmatrix}&0&\kappa&0&\kappa\\ \kappa I_{3}&0&\kappa&\kappa&\kappa\\ &0&\kappa&\kappa&0\end{pmatrix}$ [$7,3,3$]
$\begin{pmatrix}&0&\kappa&0&\kappa\\ \kappa I_{3}&\kappa&\kappa&0&0\\ &0&\kappa&\kappa&0\end{pmatrix}$ [$7,3,3$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{4}&0&\kappa&\kappa\\ &0&\kappa&\kappa\\ &0&\kappa&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{4}&0&\kappa&0\\ &\kappa&0&0\\ &\kappa&0&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{4}&\kappa&\kappa&\kappa\\ &0&\kappa&0\\ &\kappa&\kappa&\kappa\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{4}&0&0&0\\ &\kappa&\kappa&0\\ &\kappa&\kappa&0\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{4}&0&\kappa&0\\ &0&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$7,4,2$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&0\\ \kappa I_{5}&\kappa&0\\ &\kappa&0\\ &0&0\end{pmatrix}$ [$7,5,1$]
Table 8: $l_{2}$-optimal free codes (Part II).
Generator Matrix $[n,k,d]$
$\begin{pmatrix}&0&\kappa\\ &\kappa&0\\ \kappa I_{5}&0&0\\ &0&0\\ &0&0\end{pmatrix}$ [$7,5,1$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&\kappa\\ \kappa I_{5}&\kappa&\kappa\\ &\kappa&0\\ &0&0\end{pmatrix}$ [$7,5,1$]
$\begin{pmatrix}\kappa&0&0&\kappa&\kappa&0&\kappa&0\\ 0&\kappa&\kappa&\kappa&\kappa&0&0&0\end{pmatrix}$ [$8,2,4$]
$\begin{pmatrix}\kappa&0&0&\kappa&\kappa&\kappa&\kappa&\kappa\\ 0&\kappa&\kappa&\kappa&\kappa&0&0&0\end{pmatrix}$ [$8,2,4$]
$\begin{pmatrix}\kappa&0&0&0&0&\kappa&\kappa&\kappa\\ 0&\kappa&\kappa&\kappa&\kappa&0&0&0\end{pmatrix}$ [$8,2,4$]
$\begin{pmatrix}&\kappa&\kappa&0&\kappa&0\\ \kappa I_{3}&\kappa&0&\kappa&\kappa&\kappa\\ &0&\kappa&\kappa&\kappa&\kappa\end{pmatrix}$ [$8,3,4$]
$\begin{pmatrix}&\kappa&\kappa&\kappa&0\\ \kappa I_{4}&\kappa&\kappa&0&\kappa\\ &0&\kappa&\kappa&\kappa\\ &\kappa&\kappa&0&0\\ \end{pmatrix}$ [$8,4,3$]
$\begin{pmatrix}&0&\kappa&\kappa&\kappa\\ \kappa I_{4}&0&\kappa&0&\kappa\\ &\kappa&\kappa&\kappa&0\\ &0&0&\kappa&\kappa\\ \end{pmatrix}$ [$8,4,3$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &\kappa&0&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &\kappa&\kappa&0\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&\kappa\\ &\kappa&\kappa&0\\ \kappa I_{5}&\kappa&\kappa&\kappa\\ &\kappa&0&\kappa\\ &0&0&\kappa\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&\kappa\\ \kappa I_{6}&\kappa&\kappa\\ &\kappa&0\\ &0&\kappa\\ &0&\kappa\end{pmatrix}$ [$8,6,2$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&0\\ \kappa I_{6}&\kappa&0\\ &\kappa&0\\ &0&\kappa\\ &0&\kappa\end{pmatrix}$ [$8,6,2$]
$\begin{pmatrix}&0&\kappa\\ &\kappa&\kappa\\ \kappa I_{6}&0&\kappa\\ &0&\kappa\\ &0&\kappa\\ &0&\kappa\end{pmatrix}$ [$8,6,2$]
Table 9: $l_{3}$-optimal free codes.
Generator Matrix [$n,k,d$]
$\begin{pmatrix}&0&0&\kappa\\ \kappa I_{3}&\kappa&0&0\\ &0&\kappa&0\end{pmatrix}$ [$6,3,2$]
$\begin{pmatrix}&0&\kappa&\kappa&\kappa\\ \kappa I_{3}&\kappa&\kappa&0&\kappa\\ &\kappa&0&\kappa&\kappa\end{pmatrix}$ [$7,3,4$]
$\begin{pmatrix}&\kappa&\kappa&\kappa\\ \kappa I_{4}&0&\kappa&\kappa\\ &\kappa&0&\kappa\\ &\kappa&\kappa&0\end{pmatrix}$ [$7,4,3$]
$\begin{pmatrix}&\kappa&0&\kappa&\kappa&0\\ \kappa I_{3}&\kappa&\kappa&\kappa&0&0\\ &\kappa&0&\kappa&0&\kappa\end{pmatrix}$ [$8,3,4$]
$\begin{pmatrix}&\kappa&\kappa&0&0&\kappa\\ \kappa I_{3}&\kappa&0&\kappa&0&\kappa\\ &\kappa&\kappa&\kappa&0&0\end{pmatrix}$ [$8,3,4$]
$\begin{pmatrix}&\kappa&0&0&\kappa\\ \kappa I_{4}&0&\kappa&0&\kappa\\ &\kappa&\kappa&0&0\\ &\kappa&\kappa&0&\kappa\end{pmatrix}$ [$8,4,3$]
$\begin{pmatrix}&0&\kappa&0\\ &0&0&\kappa\\ \kappa I_{5}&\kappa&0&0\\ &0&\kappa&0\\ &0&\kappa&0\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&\kappa&\kappa&0\\ &0&0&\kappa\\ \kappa I_{5}&\kappa&0&0\\ &\kappa&\kappa&0\\ &0&\kappa&0\end{pmatrix}$ [$8,5,2$]
$\begin{pmatrix}&0&\kappa&0\\ &\kappa&\kappa&\kappa\\ \kappa I_{5}&0&0&\kappa\\ &\kappa&0&0\\ &\kappa&\kappa&\kappa\end{pmatrix}$ [$8,5,2$]
Table 10: $l_{4}$-optimal free codes.
Generator Matrix [$n,k,d$]
$\begin{pmatrix}&\kappa&\kappa&0&\kappa\\ \kappa I_{4}&0&\kappa&\kappa&\kappa\\ &\kappa&\kappa&\kappa&0\\ &\kappa&0&\kappa&\kappa\end{pmatrix}$ [$8,4,4$]
5 Conclusion
In this paper, we have studied hulls of free linear codes over a non-unital ring $E$. Initially, we have focused on residue and torsion codes of various hulls and found the generator matrix of the hull of a free $E$-linear code. Then, four build-up construction methods have been given for constructing free $E$-linear codes with a larger length and hull-rank from free $E$-linear codes with a smaller length and hull-rank. Some examples of codes constructed by these build-up construction methods are also given. Later, we studied the permutation equivalence of two free $E$-linear codes and then discussed the hull-variation problem. Finally, we have classified optimal free $E$-linear codes under permutation equivalence for lengths up to $8$. We would also like to point out that extending this work to other non-unital rings, as appeared in the classification of Fine Fine93 , can be an important direction for future research.
Acknowledgement
The first author gratefully acknowledges financial support from the Council of Scientific & Industrial Research (under grant No. 09/1023(16098)/2022-EMR-I), Govt. of India.
Declarations
Competing interests: All the authors confirm that there is no competing interests related to this manuscript.
Statement on Data Availability: The authors confirm that this manuscript encompasses all the data used to support the findings of this study. For any essential clarifications, requests can be directed to the corresponding author.
Use of AI tools: No Artificial Intelligence (AI) tools were utilized in the writing or preparation of this manuscript.